OpenStudy (anonymous):

What is the midpoint of a line segment with the endpoints (-6, -3) and (9, -7)? A. (1, -4.5) B. (1.5, -5) C. (-4.5, 1) D. (-5, 1.5)

OpenStudy (anonymous):

What do you think is the answer?

OpenStudy (anonymous):

what formula do i use

OpenStudy (anonymous):

You use the midpoint formula, which is \( \huge (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2}) = (x, y) \)

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

Correct!!!!